OpenStudy (anonymous):

Factor (2x-1)^3-27 as the difference of two cubes. Then simplify each factor.

OpenStudy (anonymous):

\[(2x-1)^3-3^3\]

OpenStudy (anonymous):

\[a^3-b^3 = (a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

Understood?

OpenStudy (earthcitizen):

Hint!: a=2x-1, b =3

OpenStudy (anonymous):

(2x - 1 - 3)([2x-1]^2 + 3(2x-1) + 9) and now simplify